# Cardinality of Real Numbers and Set Axiomatization

### v2

@@ 2 Discussion
Dedekind previously considered a similar problem. To visualize the function $f$, consider Dedekind’s cut $A/B$, which also defines the number $T_{y}$, with $A$ and $B$ as the disjoint components of the cut. The endpoints of the nested intervals $[a_{n},b_{n}]$ serve as elements defining the cut $A/B$, with $a_{n}$ in component $A$ and $b_{n}$ in component $B$. As $n$ increases, the gap between $a_{n}$ from $A$ and $b_{n}$ from $B$ diminishes, reducing the distance between pairs $a_{n}$ and $b_{n}$ toward zero. This progressive narrowing necessitates that the only number fitting between every pair $a_{n}$ and $b_{n}$ is $T_{y}$, the number that represents the gap created by the cut, as elucidated by Dedekind [18].
This outcome is also consistent with the Bolzano-Weierstrass (BW) and the NIP. since the intervals $[a_{n},b_{n}]$ contract to a zero-size interval, a singleton, making $T_{y}$ the only viable number within the intersection of all these intervals. Consequently, $T_{y}$ is the only number that can be inserted within all these intervals. Since the function $f$ must operate across all intervals, it necessarily generates $T_{y}$ and incorporates it into the set $S^{\prime}$. This inclusion, as it will be shown, indicates that the set $S^{\prime}$ is complete, as $f$, by generating numbers $t_{n}$, effectively bridges all gaps between algebraic numbers.
Having proven this lemma, we can now return to proving that the intersection of the sequence (7) is not empty and indeed contains the number $T_{x}$ from the set $S^{\prime}$. This conclusion directly follows from the lemma, which establishes that if the requirement is to select or insert a single number in each of the closed intervals of the sequence (5), then it is impossible to avoid inserting the number $T_{x}$ in that sequence. The sequence (7) segments share the same endpoints as the segments of sequence (5); therefore, by the lemma, if it is required to insert a single number in each segment of that sequence, and to insert numbers in all segments, it is impossible to avoid inserting $T_{x}$ since the segments in sequences (5) and (7) are identical, have the same endpoints. Consequently, if only the number $T_{x}$ can be in the intersection of the sequence (5), then only $T_{x}$ can be in the intersection of the sequence (7), as both intersections are identical, both sequences must have the same number $T_{x}$ in their intersections. It follows by the lemma that the function $f$ was compelled to insert the number $T_{x}$ in the segments of sequence (7) when acted on all segments of that sequence. To satisfy the requirement of inserting numbers in all sequence segments, the function $f$ necessarily inserted $T_{x}$, because the nested and overlapping segments dictate that only $T_{x}$ can be positioned among all of them. Thus, fulfilling the requirement to insert numbers in all sequence segments implies that the function $f$ inserted $T_{x}$ in the set $S$ when acted on all segments in the set $S$. This proves the theorem and establishes the completeness of the set $S^{\prime}$, given that an arbitrarily chosen number $T_{x}$ is present in the intersection of the sequence (7).
The AC allows the production of objects without explicitly defining them. To have a number to be included in the set $S^{\prime}$, it is not necessary to define and explicitly write the sequence representing that number. The numbers are already inserted in the set $S^{\prime}$ by the function $f$ and can be recognized or identified with AC. The AC enables recognition in the set $S^{\prime}$ of all closed intervals of the sequence representing any specific real number. Consequently, to have comprised all real numbers in the set $S^{\prime}$, it is not necessary to define all sequences of all real numbers. Instead, one can use AC to recognize all numbers in the set $S^{\prime}$, by recognizing all segments of any specific sequence and that way identifying any sequence representing any real number. Since all segments in these sequences are nonempty, the intersections of these sequences are not empty; they contain numbers defined by the sequences. Nonempty intersections of these sequences necessitate that all real numbers are in the set $S^{\prime}$. They are inserted in the set $S^{\prime}$ by the function $f$. The function $f$ is required to act on all intervals $I_{n}^{\prime}=[a_{n},b_{n}]^{\prime}$ that are in the set $S$, which also includes all intervals of the sequence (7), even if that sequence is not explicitly written because it is implicitly incorporated in the set $S^{\prime}$, as can be confirmed by AC.
The crucial part of the proof is Lemma 1, which may require more elaboration. On one hand, it is evident that in an infinite converging sequence of algebraic pairs, if one is required to either insert or select a number between each pair, the act of inserting or selecting the number defined by that specific sequence of pairs cannot be avoided since that is the only number in the intersection of all pairs, as follows from either, Dedekind’s cut, the BW theorem, or the NIP. On the other hand, one might contend that between any two algebraic pairs, it is possible to insert a number representing their average value, which would not necessarily be a transcendental number, thus seemingly obviating the need to select or insert in the sequence intervals, at any stage, the number which is defined by the sequence. However, this argument overlooks the infinite structure of the sequence, leading to a fallacy. To clarify this misconception, let us examine the infinite sequence of closed intervals $[-1/n,1/n]$.
Lemma 2: Let $\{I_{n}\}_{n=1}^{\infty}$ be an infinite sequence of nested closed intervals given as $I_{n}=\left[-\frac{1}{n},\frac{1}{n}\right]$ for $n\in\mathbb{N}$. For any function $f:\mathbb{N}\rightarrow\mathbb{R}$ assigning numbers $f(n)$ to all intervals $I_{n}$, with the constraint that $f(n)$ must lie within $I_{n}$ for all $n$, the inevitability of selecting 00 by $f$ arises not from the individual choices at any finite step, but from the collective convergence of the sequence. This is because by NIP 00 is the only point remaining within all intervals as they converge, i.e., it is the only number that satisfies $-1/n<f(n)<1/n$ for all $n$.
Proof of Lemma 2: The sequence $\{I_{n}\}_{n=1}^{\infty}$ contracts towards the center point 00 as $n$ approaches infinity. By the NIP, the intersection of this infinite sequence of nested nonempty closed intervals is non-empty, containing only the number 00. As every selected number at any step $n$ must come from the intersection of the first $n$ intervals, and 00 is the only number in the intersection of all intervals, its selection is unavoidable when going through all intervals.
This implies that the request to insert numbers in all sequence intervals requires that a single number, 00, must be in the intersection. As a result, the number 00 must be inserted in intervals while fulfilling the imposed request, ensuring the presence of numbers in all intervals. The essence of the NIP, in combination with the contracting nature of the intervals towards 00, dictates that 00 is the only number that lies within the intersection of all intervals, i.e., it is the only number within all intervals, as $n$ extends towards infinity. It is the only number that satisfies, $-1/n<0<1/n$ for all $n$. Thus, generating the number 0 with the function $f$ is unavoidable to satisfy the condition $f(n)\in I_{n}$ across the entire sequence, underscoring the inevitability of selecting 0 when considering the sequence’s collective behavior and convergence.
As explained earlier it also follows from the Dedekind’s cut $A/B$ defined by the sequence’s segments $[-1/n,1/n]$, where endpoints $-1/n$ define left component $A$ and endpoints $-1/n$ form the right component $B$. The ‘gap’ between $A$ and $B$ is precisely at 00, which separates component $A$ that contains numbers less than 00 and $B$ that contains numbers greater than 00. Since $A$ has no greatest element and $B$ has no least element (within $-1/n$ and $1/n$ numbers), the cut they form corresponds to the real number that fills this gap, which is 00.
Thus, the Dedekind cut $A/B$ demonstrates that 00 is the only number that can be situated between the two components without belonging to either, providing a unique real number that completes $\mathbb{R}$ at this point. The construction illustrates that the requirement to insert a number between every pair $-1/n$ and $1/n$ can be fulfilled only by inserting the number 00. There are no other numbers that can be between all pairs, except for the number 00.
This highlights 00 as the distinct number bridging the gap between the left and right components. This reflects the essence of Dedekind’s approach to real numbers: filling in the gaps among rational numbers to form the continuous real number line. Therefore, in the ‘gap’ created by this specific cut in the rational numbers $1/n$, 00 is identified as the only fitting number, exemplifying the completeness and continuity of $\mathbb{R}$ through Dedekind’s construction [18].
In a sequence $\left[-\frac{1}{n},\frac{1}{n}\right]$ numbers like $\frac{1}{2n}$ are absent in the intersection of all intervals, having only 00 present in the intersection of all intervals. Excluding 00 (as in $\left[-\frac{1}{n},0\right)\cup\left(0,\frac{1}{n}\right]$) leads to an empty intersection. Each interval $\left[-\frac{1}{n},0\right)$ contains $-\frac{1}{n}$, but their intersection becomes empty when all intervals are considered. Similarly, numbers like $\frac{1}{2n}$ are excluded in the limit, leaving only 00 in the intersection. Thus, the requirement to place numbers in all sequence intervals cannot be fulfilled if the number 00 is not selected, since the intersection interval will be empty.
Each interval in the sequence $\left[-\frac{1}{n},0\right)$ includes by definition the boundary point $-\frac{1}{n}$, ensuring that there are no empty intervals. This implies that the intersection of the first $n$ intervals remains nonempty, irrespective of the size of $n$. Given that each interval occupies a finite $n$-th position in the sequence and contains at least the number $-\frac{1}{n}$, one might assume that the intersection of all intervals, of the entire sequence, is also nonempty, because there is no empty intervals, each contains point $-\frac{1}{n}$. But contrarily, when considering the entire sequence, this intersection becomes empty, as all $-\frac{1}{n}$ are progressively excluded. Numbers like $-\frac{1}{n}$ do not allow to complete traversal through all members of the sequence; there will always be sequence members beyond any given number $-\frac{1}{n}$. The reason lies in the limit behavior, which differs from the behavior of individual sequence elements. Similarly, in the sequence $\left[-\frac{1}{n},\frac{1}{n}\right]$, while numbers like $\frac{1}{2n}$ fit within individual intervals, they are excluded in the collective limit, leaving only 0 present in all intervals.
In the sequence $[-1/n,1/n]$, no specific $n$-th interval mandates selecting or inserting 00 instead of $\frac{1}{2n}$. However, when considering the collective scope of all intervals, only 00 remains in their intersection, necessitating its selection. Similarly, in the sequence $[-\frac{1}{n},0)$, no single $n$ leads from a non-empty to an empty intersection. There is no specific $n$-th interval where the sequence intersection will change from non-empty to empty. It is the cumulative effect of all intervals as $n$ approaches infinity that results in this outcome, highlighting the significance of considering the entire sequence collectively. When considering all intervals $[-1/n,1/n]$ for every natural number $n$, we are essentially looking for the intersection of these intervals. The intersection of a set of nested intervals like $\left\{[-1/n,1/n]\right\}_{n=1}^{\infty}$ is the single point that is common to all intervals, and in this case, it is 00. Even when different numbers can be inserted in each interval, the requirement to insert numbers within every interval $[-1/n,1/n]$ across the entire infinite set of such intervals, and the contracting nature of these intervals to a single point, which is 00, necessitates choosing number 00 since it is the only number that fits in every interval of that infinite set.
This is an unexpected result: the requirement imposed on the function $f$ to insert a single transcendental number in each of the intervals with algebraic endpoints or between any two algebraic numbers is so compelling that it effectively fills all ‘gaps’ in the algebraic numbers, resulting in a continuum and yielding the complete set $S^{\prime}$. On the other hand, if transcendental numbers exist between every pair of algebraic numbers, then there can be no ‘gaps’. The existence of ‘gaps ’ would imply the presence of algebraic pairs without the transcendental numbers between them, which is impossible given that the function $f$ inserted transcendental numbers between every such pair of algebraic numbers.
The reason for this is that in the set $S$, confronting infinite nested sequences, such as the sequence (7), cannot be avoided, and there is a requirement that the function $f$ acts on all segments $I_{n}^{\prime}$ in these sequences. Even if the function $f$ chooses an arbitrary number from each single interval, there is no freedom which number the function $f$ will select or generate when acting on all closed nested converging intervals of an infinite sequence. The function $f$ acts on all sequence intervals and always selects or generates numbers that are in the intersection; therefore, it must select or generate the number that will be in the intersection of all intervals when it is going through all of them. Only the number defined by the sequence satisfies that requirement.
It is essential to recognize that the function $f$ operates not only on individual intervals but also on the entire collections of intervals, encompassing every conceivable sequence that can be derived from these intervals. In the set $S$, although no particular sequence is predefined, the set $S$ contains all segments necessary to construct any given sequence. The specific arrangement of these segments within $S$, whether they form a sequence or not, is irrelevant. What matters is that they are included in $S$, and that the function $f$ acts on all of them.
Utilizing the AC, one can extract within $S$ the segments pertinent to any specific sequence and organize them to represent that sequence before the function $f$ operates on them. This organization leads to an overlap of the segments for a nested sequence, guaranteeing that, in a case where segments are converging and the sequence is infinite, a unique number defined by that sequence can only be placed within all these overlapping segments. However, it is not imperative to align and overlap these segments before $f$ operates on them since the outcome should be invariant under the segments’ realignments, whether the segments are aligned and overlapping or merely present in the set unaligned, if the function $f$ must act on all of them.
The function $f$ is subjected to the same constraints in choosing numbers that fit all these segments, and the number defined by any specific sequence must be selected regardless of the segments’ arrangements. Whether these specific segments representing a sequence are aligned in the form of that sequence before $f$ acts on them, the result must be identical since $f$ will encounter all the same segments of that sequence. The definition of $f$ mandates its action on all segments in $S$, including those which are not arranged or are arranged, representing a specific sequence, with sizes diminishing towards the transcendental number defined by that sequence. Therefore, the selection of this unique number becomes inevitable, as it is the only number that can exist in all such segments, whether they are aligned or not within the sequence. There is no need to require explicit alignments since their nested structure makes theme effectively aligned.
The proof relies on the axiom of choice, which says if $TS$ is a disjointed set that does not contain the null set, the Cartesian product $PTS$ is different from the null set. In other words, the axiom maintains that its assumptions fulfilled, among the subsets of $UTS$ (or selection-set of TS) there is at least one whose intersection with each member of $TS$ is a unit-set. This ensures that after selecting all intervals $[a_{n},b_{n}]^{\prime\prime}$ in the set $S^{\prime}$, which are the same as the intervals $[\alpha_{n},\beta_{n}]$ in the sequence representing the missing number $T_{x}$, the intersection of these intervals will contain the single number $T_{x}$.
This unexpected result and its significance necessitate an independent proof of the completeness of the set $S^{\prime}$, which is provided by the following theorem.
Theorem 2: Define $S^{\prime}$ as the set comprising all algebraic numbers and only transcendental numbers uniquely determined by the function $f:S\rightarrow\mathbb{R}\setminus\overline{\mathbb{Q}}$, where $S$ includes every algebraic interval $I_{n}^{\prime}=[a_{n},b_{n}]\cap\overline{\mathbb{Q}}$ with algebraic endpoints satisfying $a_{n}\leqslant b_{n}$, and each interval $I_{n}^{\prime}$ is defined such that it contains only algebraic numbers. For each such interval $I_{n}^{\prime}$, the function $f$ assigns a transcendental number $f(I_{n}^{\prime})=t_{n}$, ensuring $t_{n}$ falls within $I_{n}^{\prime}$ and belongs to the set $\mathbb{R}\setminus\overline{\mathbb{Q}}$, i.e., $t_{n}$ is not algebraic, and the function $f$ has to act on all intervals $I_{n}^{\prime}$ in the set $S$. The set $S^{\prime}$ constructed in this manner is everywhere dense in the sense of Dedekind, meaning it has no ‘holes’. Specifically, for every open interval within $\mathbb{R}$, there exists at least one element $t_{n}$ of the set $S^{\prime}$ within each such interval. These $t_{n}$ elements act as ‘bridges’ between algebraic numbers, covering all possible transcendental ‘holes’ in the set $S^{\prime}$.
Proof of Theorem 2: The proof is straightforward since if between every two algebraic numbers there is a transcendental number $t_{n}$, then pairs of algebraic numbers in a Dedekind cut cannot create a ‘hole’ in that set. A ‘hole’ would require that there are algebraic pairs defining the cut between which there is no $t_{n}$ transcendental number. However, let us formalize this proof. To prove the theorem, it will be shown that any cut $A/B$ in the set $S^{\prime}$ does not create a ‘hole’ in $S^{\prime}$ and therefore does not generate new elements in $S^{\prime}$ since the first component $A$ will have the last element or the second component $B$ will have the first element, or both conditions will occur.
The cut $A/B$ will also make a cut in the set of algebraic numbers $\overline{\mathbb{Q}}$ in $S^{\prime}$. This cut defines a well-defined number
$t=(A\cap\overline{\mathbb{Q}})/(B\cap\overline{\mathbb{Q}})$ (9)
The same cut $A/B$ splits intervals $I_{n}^{\prime\prime}=[a_{n},b_{n}]\cap\overline{\mathbb{Q}}\cup T$ with algebraic endpoints $a_{n},b_{n}$ in the set $S^{\prime}$, positioning all endpoints $a_{n}$ within the cut component $A$, and all endpoints $b_{n}$ within component $B$. This configuration leads to an infinite nested sequence of intervals $I_{n}^{\prime\prime}$, facilitated by the everywhere dense nature of $\overline{\mathbb{Q}}$, with the sizes of the intervals contracting towards zero. The number $t$ is similarly defined by a sequence of these nested intervals with algebraic endpoints $a_{n},b_{n}$
$[a_{1},b_{1}]^{\prime\prime}\supseteq[a_{2},b_{2}]^{\prime\prime}\supseteq[a_{3},b_{3}]^{\prime\prime}...[a_{n},b_{n}]^{\prime\prime}\supseteq[a_{n+1},b_{n+1}]^{\prime\prime}\supseteq...$ (10)
Both numbers $t$, defined by the cut (9) in algebraic numbers, and $t$, defined by the sequence (10) of intervals $I_{n}^{\prime\prime}$, are the same because they are defined by the same cut $A/B$ in $\overline{\mathbb{Q}}$ and therefore must be identical.
Since all algebraic intervals of the sequence (10) are nested and nonempty, because the function $f$ by definition inserts a transcendental number in each of them, they will contain numbers common to all intervals, as required by their nested structure. The same numbers appear in all nested intervals $[a_{n},b_{n}]^{\prime\prime}$ because $[a_{(n+1)},b_{(n+1)}]^{\prime\prime}\subseteq[a_{n},b_{n}]^{\prime\prime}$ for every $n$. In these nested intervals, any new point added by the construction in the interval $[a_{(n+1)},b_{(n+1)}]^{\prime}$ will also appear in the preceding intervals $[a_{n},b_{n}]^{\prime}$, $[a_{(n-1)},b_{(n-1)}]^{\prime}$, $[a_{(n-2)},b_{(n-2)}]^{\prime}$, …, $[a_{1},b_{1}]^{\prime}$. Consequently, there will be an intersection (an interval) that will contain these common numbers from the overlapping intervals. Because the sequence is infinite and the intervals are contracting, the interval representing the intersection of all intervals reduces to a zero-size interval with only one number remaining, the number $t$, defined by the sequence (10).
The number in the intersection, by definition, is one of the numbers from the intervals of the sequence (10). Therefore, it is one of the $t_{n}$ numbers created by the function $f$ when it acted on the intervals $[a_{n},b_{n}]^{\prime}$, since only the $t_{n}$ numbers are present in the segments of the sequence (10). Because the numbers generated by the function $f$ are elements of the set $S^{\prime}$, the number $t$ must be from either component $A$ or component $B$ of the set $S^{\prime}$.
It will be shown that if $t\in A$, then it is the last element of component $A$.
@@
If $t$ is not the last element of component $A$, then there is
@@
The $t^{\prime}$ cannot be an algebraic number because $t<t^{\prime}$ and (9), which states that $t$ is a cut, will require that $t^{\prime}\in(B\cap\overline{\mathbb{Q}})$, and therefore $t^{\prime}\in B$, which contradicts that $t^{\prime}\in A$. If $t^{\prime}$ is a transcendental number, then it represents a cut $C/D$ in the set of algebraic numbers $\overline{\mathbb{Q}}$, which creates a hole in $\overline{\mathbb{Q}}$, and for this reason, the first component $C$ of that cut $t^{\prime}$ does not have the last element.
Because $t<t^{\prime}$ implies $(A\cap\overline{\mathbb{Q}})/(B\cap\overline{\mathbb{Q}})<C/D$, it follows that $(A\cap\overline{\mathbb{Q}})\subset C$. If $t^{\prime\prime}$ is any number from the set $C\setminus A$ (because the set $S^{\prime}$ is everywhere dense), then $t<t^{\prime\prime}$ requires $t^{\prime\prime}$ to be in $B$ (since $t^{\prime\prime}\in S^{\prime}\setminus A=B$).
@@
Contrarily, because $t^{\prime\prime}\leqslant t^{\prime}$ (as $t^{\prime}$ represents a cut in $\overline{\mathbb{Q}}$ and is in $A$, as referenced in (11)), it requires that $t^{\prime\prime}\in A$. However, having $t^{\prime\prime}$ in both $A$ and $B$ is not possible because $A$ and $B$ are disjoint components of the cut. Therefore, there can be no $t^{\prime}\in A$ such that $t<t^{\prime}$.
Thus, if $t\in A$, then it must be the last element of component $A$. Similarly, if $t\in B$, it can be shown in the same manner that $t$ is the first element of the second component $B$.
Therefore, any cut in the set $S^{\prime}$ does not produce a hole or generate new elements, which substantiates Theorem 2.
Explicit sequence of real numbers

### v3

@@ 2.1 Main Construction and Theorem
Dedekind considered a similar problem. The number $r$ can be defined through Dedekind’s cut $A\mid B$, with $A$ and $B$ as the disjoint components of the cut. The endpoints of the nested intervals $[a_{n},b_{n}]$ serve as elements defining the cut $A\mid B$, with $a_{n}$ in component $A$ and $b_{n}$ in component $B$. As $n$ increases, the gap between $a_{n}$ from $A$ and $b_{n}$ from $B$ diminishes, reducing the distance between pairs $a_{n}$ and $b_{n}$ toward zero. This progressive narrowing necessitates that the only number fitting between every pair $a_{n}$ and $b_{n}$ is $r$, the number that represents the gap created by the cut, as it was shown by Dedekind [16].
Since Theorem 1. is an important result, let us derive it independently.
3 Alternative proof for Completeness of the set $S_{m}$
Theorem 2 (Dedekind Completeness of $S_{m}$ via Dedekind Cuts).
The set $S_{m}$ as defined in Theorem 1, is Dedekind complete. Any Dedekind cut $A\mid B$ of $S_{m}$ into two disjoint components $A$ and $B$ does not create a gap. Specifically, the first component $A$ has a greatest element, or the second component $B$ has a least element, or both. This means the set $S_{m}$ as defined in Theorem 1, contains all real numbers, there is no a number that is not marked. Any new different marking will not produce marked numbers that are not already included in the marked set $S_{m}$.
Proof.
Let $S_{Q}$ and $S_{m}$ be, as defined in theorem 1. The set $S_{Q}$ consists of all closed intervals $[a_{\alpha},b_{\alpha}]$ with rational endpoints $a_{\alpha},b_{\alpha}\in\mathbb{Q}$, $a_{\alpha}\leq b_{\alpha}$. The set $S_{m}$ is constructed by marking one irrational number $x_{\alpha}$ in each interval $[a_{\alpha},b_{\alpha}]\in S_{Q}$, under the conditions specified in the theorem, and includes all such numbers as well as all rational numbers $\mathbb{Q}$.
Relationship Between Cuts in $\mathbb{Q}$ and $S_{m}$
Marking an irrational number $r\in S_{m}$ in every interval $[a_{\alpha},b_{\alpha}]\in S_{Q}$ corresponds to creating Dedekind cuts $A\mid B$ in $\mathbb{Q}$. Specifically, each marked $r$, divides the rational numbers $\mathbb{Q}$ into two disjoint subsets: $A=\{q\in\mathbb{Q}:q<r\},\quad B=\{q\in\mathbb{Q}:q\geq r\}.$ This partition satisfies the properties of a Dedekind cut: $A\cup B=\mathbb{Q}$, $A\cap B=\emptyset$, and for all $a\in A$, $b\in B$, $a<b$.
So, marking an irrational number in an $S_{Q}$ interval is equivalent to representing a gap in $\mathbb{Q}$ within that interval.
Demonstrating Completeness of $S_{m}$
After all intervals $I_{n}$ are marked and set $S_{m}$ is composed, consider inspecting that all irrational numbers are marked by performing an additional, new arbitrary marking of already marked intervals $I_{n}$ and check if such additional markig will producen a different marked number $t$ that is already not icluded in the marked set $S_{m}$. Such additional marking will split all numbers in the interval, including marked numbers, into two disjoint components $A$ and $B$. So, it represents a Dedekind cut
$A\mid B$
in an interval $I_{n}$ and also a cut in the set of marked numbers $S_{m}$ separating them also in the components $A$ and $B$.
Since any marked number by the theorem is irrational, the new marked number $t$ also makes a Dedekind cut in rational numbers $\mathbb{Q}$:
$t=(A\cap\mathbb{Q})\mid(B\cap\mathbb{Q}).$ (13)
As a marked number $t$ belongs to the set of marked numbers $S_{m}$, and as a marked element of the interval $I_{n}$, it must belong to one of the components $A$ or $B$ of the cut in $I_{n}$, therefore it belongs to either of the components $A$ or $B$ of both $S_{m}$, and $I_{n}$.
It will be shown that this new marking does not create a gap in the set $S_{m}$ of marked numbers. Specifically, the first component $A$ of marked numbers has a greatest element, or the second component $B$ has a least element, or both. So, it will be shown that any new marking generates a marked numbers that already exist in set $S_{m}$.
@@ 3 Alternative proof for Completeness of the set $S_{m}$
Let us show that if $t\in A$, then it is the greatest element of component $A$. Suppose $t$ is not the greatest element of component $A$. Then there exists a
@@
The $t^{\prime}$ cannot be a rational number because $t<t^{\prime}$ and, by (13) the definition of the cut $t$, it would require $t^{\prime}\in B\cap\mathbb{Q}$, contradicting $t^{\prime}\in A$.
If $t^{\prime}$ is an irrational number, it represents a cut $C\mid D$ in $\mathbb{Q}$, which creates a gap in $\mathbb{Q}$, and for this reason, the first component $C$ of that cut $t^{\prime}$ does not have the greatest element. Because $t<t^{\prime}$, we have $(A\mid B)<C\mid D$, implying $(A\cap S_{m})\subset C$. If $t^{\prime\prime}$ is any marked number from $C\setminus A$ (since $C$ does not have the greatest element and $S_{m}$ is everywhere dense), then $t<t^{\prime\prime}$ would require $t^{\prime\prime}$ to be in $B$. Contrarily, because $t^{\prime\prime}\leqslant t^{\prime}$ (as $t^{\prime}$ represents a cut in $\mathbb{Q}$ and is in $A$, as referenced in (14)), it requires that $t^{\prime\prime}\in A$.
@@
However, having $t^{\prime\prime}$ in both $A$ and $B$ is not possible because $A$ and $B$ are disjoint components of the cut. This contradiction implies there is no $t^{\prime}\in A$ such that $t<t^{\prime}$, and so $t$ must be the greatest element of $A$.
Similarly, if $t\in B$, it can be shown that $t$ is the least element of component $B$.
This ensures that $S_{m}$ satisfies Dedekind completeness: any cut in $S_{m}$ does not create a gap or require additional elements. Since $t$ is either the greatest element of $A$ or the least element of $B$, it follows that any Dedekind cut $A\mid B$ in $S_{m}$ does not create a gap. Every cut is represented by a unique marked point $t\in S_{m}$, ensuring that $S_{m}$ satisfies Dedekind completeness.
Thus, $S_{m}$ is complete in the Dedekind sense, having no gaps, affirming the Theorem. ∎
Let us show Dedekind complentess also on this independet way:
Theorem 3 (Dedekind Completeness of $S_{m}$ via Missing Elements).
The set $S_{m}$, as defined in Theorem 1, is Dedekind complete. That is, $S_{m}$ contains every real number: no real number is missing from $S_{m}$.
Proof.
Suppose, for contradiction, that there exists a real number $t_{x}\notin S_{m}$. Then $t_{x}$ determines a Dedekind cut in $\mathbb{Q}$, and can be represented as the unique limit of a nested sequence of closed intervals with rational endpoints:
$I_{n}=[a_{n},b_{n}],\quad a_{n},b_{n}\in\mathbb{Q},\quad I_{1}\supseteq I_{2}\supseteq\cdots,\quad\bigcap_{n=1}^{\infty}I_{n}=\{t_{x}\}.$
Since $S_{Q}$ enumerates all intervals with rational endpoints, each $I_{n}\in S_{Q}$, and the sequence $\{I_{n}\}$ is a subsequence of this enumeration. By the global, simultaneous marking condition (S3), every interval and all its subintervals are assigned a mark from the outset.
Thus, for each $n$, the interval $I_{n}$ contains not just one marked point but infinitely many, inherited through the structural marking:
$M\cap I_{n}\neq\emptyset\quad\text{for all }n\in\mathbb{N}.$
Since the sequence $\{I_{n}\}$ is nested and every $I_{n}$ intersects $M$, the finite intersections remain nonempty:
$\bigcap_{n=1}^{m}(M\cap I_{n})\neq\emptyset\quad\text{for all }m\in\mathbb{N}.$
Now consider the infinite intersection. Because the intervals satisfy $\bigcap_{n=1}^{\infty}I_{n}=\{t_{x}\}$, we obtain:
$\bigcap_{n=1}^{\infty}(M\cap I_{n})=M\cap\left(\bigcap_{n=1}^{\infty}I_{n}\right)=M\cap\{t_{x}\}.$
To conclude $t_{x}\in M$, we reason as follows:
Every interval $I_{n}$ contains both the point $t_{x}\in I_{n}$ and at least one marked point $x_{n}\in M\cap I_{n}$. By condition (S3), the marking is structural and simultaneous, so no “gaps” are left that would prevent progression through the entire nested sequence.
Since $t_{x}$ survives through all $I_{n}$, and some mark exists in each interval together with $t_{x}$, structural marking ensures that at least one of these marks also persists through all intervals, i.e., there exists some $x_{k}\in M$ that lies in every $I_{n}$. All other marks $x_{n}\neq t_{x}$ will be eliminated with smaller intervals. Thus to avoid leaving a deep interval unmarked, the limit point must have been one of marks,
But the only point that lies in every $I_{n}$ is $t_{x}$. Therefore, $x_{k}=t_{x}$, and so $t_{x}\in M\subseteq S_{m}$, contradicting our assumption.
Thus, $t_{x}$ is a marked point, and every real number lies in $S_{m}$, proving Dedekind completeness. ∎
4 Explicit sequence of everywhere dense set of real numbers